mercvriii:

In standard form, write an equation for the line that passes through (–2, –5) and is parallel to the line 15x − 5y = 10.

jimthompson5910:

Anything parallel to Ax+By = C will be of the form Ax+By = D The only difference is the C has been replaced with D. For example, 2x+3y = 5 is parallel to 2x+3y = 7 Both equations have a slope of -2/3. Parallel lines have equal slopes but different y intercepts.

jimthompson5910:

In the case of 15x − 5y = 10, we see that A = 15, B = -5 and C = 10 Anything parallel to this will be 15x - 5y = D To find the value of D, plug in (x,y) = (-2, -5), which is the point we want the parallel line to go through. Simplify to calculate D.

mercvriii:

@jimthompson5910 Thank you, this helped! :)

jimthompson5910:

You're welcome. Let me know what you get for that D value.

mercvriii:

@jimthompson5910 I got −5≠10, which is false.

jimthompson5910:

No, it's a good thing you didn't get 10 as a result. This is because the y intercepts are different for the parallel equations. D = -5 is correct. The equation 15x - 5y = -5 is parallel to 15x-5y = 10 The point (-2, -5) is on 15x - 5y = -5

jimthompson5910:

If you wanted, you can solve each for y 15x - 5y = -5 -5y = -15x - 5 y = (-15x-5)/(-5) y = (-15x)/(-5) - 5/(-5) y = 3x + 1 Similarly 15x - 5y = 10 -5y = -15x + 10 y = (-15x+10)/(-5) y = (-15x)/(-5) +10/(-5) y = 3x - 2 Both equations have a slope of m = 3, but different y intercepts (b = 1 and b = -2 respectively)

mercvriii:

@jimthompson5910 I was confused with this question for hours, thank you. <3

jimthompson5910:

You're welcome